OpenStudy (anonymous):

What is the simplified expression for 3 to the power of negative 4 multiplied by 2 to the power of 3 multiplied by 3 to the power of 2 whole over 2 to the power of 4 multiplied by 3 to the power of negative 3? 3 over 2 3 to the power of 2 over 2 to the power of 2 3 to the power of 2 over 2 2 to the power of 4 over 3

OpenStudy (anonymous):

PLEASE HELP ME!!!!

OpenStudy (amoodarya):

\[\frac{3^{-4}*2^{3}*3^2}{2^4*3^{-3}}\] is it ?

OpenStudy (anonymous):

yes! i tried so hard do you know the answer

OpenStudy (amoodarya):

\[\frac{3^{-4}*3^2*2^{3}}{2^4*3^{-3}}=\\\frac{3^{-4}*3^2*3^{+3}}{2^4*2^{-3}}=\\\frac{3^{-4+3+3}}{2^{4-3}}\]

OpenStudy (amoodarya):

\[a^{m}=\frac{1}{a^{-m}}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what do you think the answer is

OpenStudy (amoodarya):

\[\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

THANK YOU!!

OpenStudy (amoodarya):

your welcome but ,do you understand it ?